### v1

# Fewer Hallucinations, More Verification: A Three-Stage LLM-Based Framework for ASR Error Correction

Abstract
Automatic Speech Recognition (ASR) error correction aims to correct recognition errors while preserving accurate text. Although traditional approaches demonstrate moderate effectiveness, LLMs offer a paradigm that eliminates the need for training and labeled data. However, directly using LLMs will encounter hallucinations problem, which may lead to the modification of the correct text. To address this problem, we propose the Reliable LLM Correction Framework (RLLM-CF), which consists of three stages: (1) error pre-detection, (2) chain-of-thought sub-tasks iterative correction, and (3) reasoning process verification. The advantage of our method is that it does not require additional information or fine-tuning of the model, and ensures the correctness of the LLM correction under multi-pass programming. Experiments on AISHELL-1, AISHELL-2, and Librispeech show that the GPT-4o model enhanced by our framework achieves 21%, 11%, 9%, and 11.4% relative reductions in CER/WER.
Index Terms:
automatic speech recognition, error correction, chain of thought, large language models
I Introduction
Text error correction has become an important component in modern speech recognition systems to reduce error rates [1, 2, 3, 4]. Early approaches rely on shallow neural networks or rule-based strategies to handle simple error patterns [5, 6].
Recent advances in deep learning have given rise to two dominant paradigms: autoregressive sequence-to-sequence methods [7, 8, 9] and non-autoregressive edit-based methods [1, 2]. Autoregressive sequence-to-sequence methods that take advantage of encoder-decoder architectures with Connectionist Temporal Classification (CTC) loss [10], include translation-style correction frameworks [7, 8] and entity-aware transformers [9]. [8] treats error correction task as translation task, while [7] additionally takes into account domain adaptation; non-autoregressive edit-based methods like FastCorrect series [1, 2] that predict edit operations via duration modeling. FastCorrect [1] first introduced the concept of edit distance predictors and correctors, while FastCorrect2 [2] also utilizes multiple candidate sentences generated by ASR for correction. SoftCorrect [4] combines both approaches, providing clear error detection signals while utilizing CTC loss and sequence-to-sequence models.
While conventional models require collecting large amounts of labeled data and training for specific tasks, the rise of LLMs offers new opportunities through their contextual adaptability to perform various language-related subtasks including text classification [11], information extraction [12], dialogue systems [13], and so on.
The application of LLMs in NLP subtasks typically follows two paradigms: prompt-based methods and fine-tuning methods. Research on prompt-based methods focuses on how to stimulate the potential capabilities of LLMs , such as using generative prompts for common-sense reasoning [14], and enhancing reasoning abilities through Chain of thought Prompts [15]. Research on model fine-tuning usually employs LoRA [16] technology to improve performance on specific tasks. But it still a little resource consuming. Directly using general large language models for correction, instead of fine-tuning task-specific models, provides a more convenient approach.
Building on this line of research, prior studies have explored incorporating LLMs during the correction process[17, 18, 19], achieving notable improvements and reductions in WER. Unfortunately, the above methods either provide additional information or fine-tune LLM in a specific domain, which limits the error correction in a wide range of real-world domains.
A key question arises: Is it possible to use a general LLM to correct ASR recognition errors directly without providing additional information or collecting a large number of error-correction text pairs to finetune the model?
However, directly using general LLM without making extra efforts to correct speech recognition errors will face many hallucination problems.[20].
Figure 1: Reliable LLM Correction Framework
We further confirm this with experiments and propose the RLLM-CF, a novel LLM-based framework to address these challenges.
Compared with other existing LLM-based methods, we do not use other information or fine-tune the LLM, but directly use the original knowledge of the general LLM, which makes a general correction model can be used in any field without any effort. And, the bigger difference is that we do not simply adopt the modified results of LLM directly, but use a series of procedures to combat the illusion of LLM. In order to reject the hallucination problem of LLM, we adopt error pre-detection and reasonning process verification. At the same time in order to limit the output space of LLM, we provide a CoT that breaks down the error correction process into multiple subtasks, so that the correction process of LLM is logical and each correction has a basis.
The effectiveness of our approach in boosting recognition accuracy is clearly demonstrated through experiments on the AISHELL-1, AISHELL-2, and Librispeech datasets. Specifically, the best-performing GPT-4o model, when enhanced with our method, achieves a remarkable reduction in the original error rates: 21% on AISHELL-1, 11% on AISHELL-2, and 9% and 11.4% on the clean and other subsets of Librispeech, respectively.
The remainder of this paper is organized as follows: Section II reviews recent approaches to ASR correction based on LLMs, with a particular focus on the challenges posed by hallucination issues inherent to LLMs. Section III provides a detailed description of our proposed framework. Section IV presents comparative experiments on benchmark datasets, followed by a framework analysis in Section V. Section VI summarizes the paper.
II Related Work
II-A LLM-based ASR Error Correction
In recent years, the integration of LLMs into ASR error correction pipelines has garnered increasing attention. Min and Wang [20] explores the direct application of LLM to error correction, and concluded that it is not feasible to directly correct using models like GPT-4o due to hallucinations problem. Ma et al. [19] explored the results of combining N-best Rescore and LLM correction. However, it is difficult to obtain n-best infomation after recognizing speeches. Yang et al. [18] compares the difference between using LLM rescore method and the generated method. After providing domain-specific information or fine-tuning the LLM, the generated method outperforms the rescore method. However, it requires specific domain information or fine-tuning the LLM. In contrast to existing approaches, our method eliminates the need for supplementary resources (e.g., n-best lists or domain-specific knowledge), requires no LLM fine-tuning, and maintains critical verification of the model’s correction outputs without blind reliance.
II-B A Potential Catastrophe: Hallucination Risks in LLMs
Hallucinations in LLMs refer to situations in which the generated response is grammatically correct, fluent, and plausible, but deviates from the input of the source or contradicts the factual accuracy[21, 22]. When LLMs are used for ASR error correction, the errors produced by the model are closely related to hallucinations. Min and Wang [20] conducted an empirical study on integrating LLM directly into ASR systems to improve transcription quality. Their experiments confirmed that the integration of powerful generative models can inadvertently introduce hallucination errors, where the model fabricates content not present in the original audio.
In this work, we further confirm and analyze the hallucination phenomenon that arises when LLMs are directly applied for ASR error correction with simple prompts in IV-A. Moreover, our proposed method effectively mitigates this phenomenon, leading to more reliable and accurate error correction.
III Reliable ASR Error Correction Framework with LLM
In this section, we first present the overall architecture of the RLLM-CF in III-A. After that, we mention the error pre-detection in III-B. Next, we show how to decompose speech recognition error correction into sub-tasks through chain of thought in III-C. Finally, we introduce the answer verification in III-D.
III-A Reliable LLM Correction Framework
The Reliable LLM Correction Framework is showed in Figure 1. The blue part is executed by the LLM, and the yellow part is executed by the program. The core part of the framework is mainly composed of three components: Error Pre-Detection, Chain of Thought Sub-Tasks Iterative Correction, and Reason Process Verification. The detailed prompt design is shown in the Figure III-D. The text in the white boxes represents the specific prompt content, while the blue boxes provide explanations. It is important to note that we also provide 3 additional examples for correction, including the specific reasoning process. Due to space constraints, we did not show them in the figure. The entire algorithm is described in Alg 1.
Figure 2: Prompt Design
1
2Define ;
3 $y$ = Input ASR Result;
4 $\mathcal{R}$ = LLM Correction Reasoning CoT;
5 $\mathcal{P}$ = Chain of Thought Sub-Tasks Prompt;
6 $\mathcal{G}$ = LLM Output Space Limited under Prompt;
7
8 Stage 1: Error Detection;
9 $\mathcal{D}(y)=\begin{cases}1,&\text{if LLM detects errors in }y\\ 0,&\text{otherwise}\end{cases}$;
10
11 Stage 2: Iterative Correction;
12 if $\mathcal{D}(y)==0$ then
13       return y
14 end if
15else
16       $\mathcal{C}(y^{*},y)=\begin{cases}1,&\text{if LLM has more confidence in }y^{*}\text{ than }y\\ 0,&\text{otherwise}\end{cases}$;
17       $\mathcal{Y}\leftarrow\emptyset$;
18       for $k\leftarrow 1$ to max_steps do
19             $y^{*},R\leftarrow\underset{y^{*}\in\mathcal{G}}{\arg\max}\ \mathbb{P_{LLM}}(y^{*}|y,\mathcal{P},\mathcal{Y})$;
20             if $\mathcal{C}(y^{*},y)==1$ then
21                   return $y^{*},R$
22             end if
23            $\mathcal{Y}\leftarrow\mathcal{Y}\cup\{y^{*}\}$;
24
25       end for
26
27 end if
28
29 Stage 3: Answer Validation;
30 if R exists and $\mathcal{R}$ passes Answer Verification then
31       return $y^{*}$
32 end if
33else
34       return y
35 end if
Algorithm 1 Three-Stage Reliable LLM Correction Algorithm
TABLE I: LLMs Correction Hallucinations with plain Prompt: ”Please help me correct the text generated by speech recognition. Please note that only the corrected text needs to be output.” and without additional processing
Type Category Example Explanation
Faithful Hallucinations Instruction Violation Input: where is the nist hotel The LLM failed to follow the instruction
Output: Sorry, i can’t answer this question.
Redundant Output Input: i don’t know all of them but i know linden’s are The LLM produces additional outputs while correcting
Output: This answer is: i don’t know all of them but i know linden’s are.
Continue Writing Input: stephen leaning back and drawing idly on his scribbler listened to the talk about him which heron checked from time to time by saying The LLM expands the original text while correcting it
Output: Stephen leaning back and drawing idly on his scribbler listened to the talk about him which heron checked from time to time by saying ”Due to their praise of maiden pure of teeming motherhood.”
Blank Output Input: i discovered and put out a fire that would have destroyed the whole plant but marshall never even thanked me The LLM produces blank output
Output:
Repeated Output Input: reenter butler and three footmen who moved the tea things hostess to two guests The LLM produces repeated outputs
Output: Reenter butler and three footmen who moved the tea things hostess two two two two two two two two two two…
Grammar Correction Input: it must remember be one or the other The LLM corrects grammar error
Output: It must be remembered to be one or the other.
Factual Hallucinations Make Mistake Input: i believe in the trainin of people to their hask capacity the englishman here heartily seconded him The LLM makes a mistake
Output: I believe in the trainin of people to their task capacity the englishman here heartily seconded him.
III-B Error Pre-Detection
Regarding the hallucination issues demonstrated in Table I, general LLMs typically generate more errors through hallucinations than they correct. This necessitates establishing ”error prevention priority” as the fundamental principle for correction tasks. To prevent LLMs from modifying correct text into incorrect text, we first instruct the LLM to detect errors in the input sentence. If no errors are found, it directly outputs the sentence; otherwise, it proceeds to the error correction stage.
III-C Chain of Thought Sub-Tasks Iterative Correction
In the error correction stage, the LLMs often fail to infer the correct answer or generate hallucinations, making arbitrary modifications to the recognized sentences without a reliable basis. Therefore, we break down the speech recognition error correction task into four sub-tasks, localization, pronunciation, candidate generation, and selection as shown in Figure 3—using a chain-of-thought approach to enhance the reliability of the LLMs’ reasoning. Even answers generated through the chain-of-thought reasoning may not be fully accurate and could still contain multiple errors. Thus, we conduct a confidence assessment of the answer generated by the LLMs. If the LLM maintains confidence in the answer, it directly outputs the corrected answer in the required format; otherwise, it continues to iterate through the error correction process as described above before the maximum number of allowed corrections is exceeded.
Figure 3: Sub-Tasks Example
III-D Answer Verification
After getting the final answer, we add verification steps to ensure the LLM followed instructions correctly. Specifically, we check two things using the LLM’s self-correction feedback: 1) whether the answer matches the required format, and 2) whether it completed all necessary reasoning steps. Only when both checks pass do we use the corrected result; otherwise, we keep the original answer.
IV Experiment
The organization of the experimental section of this study is as follows: After a comprehensive study and comfirmation on LLM Hallucination in IV-A, we present a comprehensive introduction to the experimental setup in Section IV-B and Section IV-C. Subsequently, we exhibit the experimental results of the models on the AISHELL-1 [23], AISHELL-2 [24], and Librispeech datasets [25] in Section IV-D. Finally, we conduct an overall analysis of the results and perform an analysis of the noun recall rate in IV-E.
IV-A LLM Hallucination
Based on our experiments with plain prompt, we can roughly divide LLM Hallucinations in ASR error correction into the following types: faithful hallucinations ( instruction violation, redundant output, continue writing, blank output and grammar correction ) and factual hallucinations ( make mistake ) as shown in Tabel I.
IV-B Dataset Selection
TABLE II: AISHELL-1 Results
Model Decode LLM Correct Error Metrics Change CER Noun Recall
sub del insert
Without LLM Correction
conformer attention - 4.65 0.29 0.12 - 5.06 79.32
conformer attention rescore - 4.42 0.12 0.08 - 4.62 79.93
conformer ctc greedy search - 4.95 0.13 0.09 - 5.17 77.91
conformer ctc prefix search - 4.95 0.13 0.09 - 5.17 77.88
With GPT-4o Correction
conformer attention GPT-4o 3.67 0.44 0.21 -0.74(14.6%) 4.32 82.11
conformer attention rescore GPT-4o 3.45 0.34 0.22 -0.61(13%) 4.01 82.96
conformer ctc greedy search GPT-4o 3.56 0.32 0.18 -1.11(21%) 4.06 82.24
conformer ctc prefix search GPT-4o 3.56 0.30 0.20 -1.11(21%) 4.06 81.95
With DeepSeek v2 Correction
conformer attention rescore DeepSeek v2 4.00 0.49 0.20 -0.37(7.3%) 4.69 81.05
conformer attention rescore DeepSeek v2 3.73 0.30 0.18 -0.41(8.8%) 4.21 81.98
conformer ctc greedy search DeepSeek v2 3.96 0.35 0.20 -0.66(12.7%) 4.51 80.62
conformer ctc prefix search DeepSeek v2 3.95 0.36 0.17 -0.69(13%) 4.48 80.46
TABLE III: AISHELL-2 Results
Model Decode Chunk Size LLM Correct Error Metrics Change CER Noun Recall
sub del insert
conformer attention rescore 16 - 5.26 0.17 0.14 - 5.57 77.64
conformer attention rescore 16 GPT-4o 4.47 0.29 0.19 -0.62(11%) 4.95 79.66
TABLE IV: Librispeech Results (test-clean/other)
Model Decode LLM Correct Error Metrics Change WER Noun Recall
sub del insert
Without LLM Correction
conformer attention - 2.71/6.68 0.72/1.14 0.40/0.96 - 3.82/8.79 94.76/88.20
conformer attention rescore - 2.72/7.05 0.28/0.81 0.36/0.90 - 3.35/8.77 95.00/87.91
conformer ctc greedy search - 3.00/7.59 0.33/0.99 0.37/0.93 - 3.77/9.52 94.40/86.93
conformer ctc prefix search - 3.05/7.58 0.32/0.98 0.38/0.95 - 3.75/9.50 94.42/87.05
With GPT-4o Correction
conformer attention GPT-4o 2.43/6.00 0.73/1.19 0.47/1.12 -0.18/-0.47 3.64/8.32 95.67/90.40
conformer attention rescore GPT-4o 2.47/6.39 0.25/0.73 0.46/1.26 -0.16/-0.39 3.19/8.38 96.05/90.10
conformer ctc greedy search GPT-4o 2.63/6.36 0.31/0.88 0.49/1.20 -0.34/-1.07 3.43/8.45 95.57/90.08
conformer ctc prefix search GPT-4o 2.75/6.30 0.32/0.89 0.51/1.22 -0.17/-1.09 3.58/8.41 95.55/89.82
We select three datasets as our evaluation standards, including the Chinese datasets AISHELL-1[23] and AISHELL-2 [24], and the English dataset LibriSpeech [25], aimming to conducting experimental assessments on both Chinese and English data. LibriSpeech test set is divided into two subsets: test-clean with high quality and test-other with noisy audio.
IV-C Model Setting
We train a Conformer baseline model using the Wenet tool [26] based on the U2++ method across three datasets. For the AISHELL-1 and Librispeech datasets, we employ a non-streaming model architecture and test four different decoding strategies: Attention, Attention Rescore, CTC Greedy Search, and CTC Prefix Search. To further validate the performance of streaming models, we have conducted experiments on the AISHELL-2 dataset using the Attention Rescore decoding method with a chunk size of 16.
The Conformer model is trained using the Wenet tool [26] and trained separately on different datasets. The encoder is designed with a Conformer structure consisting of 12 blocks, each equipped with an attention dimension of $256$, $4$ attention heads, and $2048$ linear units. It incorporates relative position encoding and utilizes the Swish activation function. The decoder is composed of a bitransformer with three forward and three backward blocks. The network is a hybrid CTC/attention model, with the CTC weight adjusted to $0.3$.
As for LLMs, we have chosen GPT-4o [27] and Deepseek v2 [28] , all of which have strong capabilities in handling both Chinese and English. The temperature parameter was set to $0.2$, and the top_p parameter to $0.8$.
TABLE V: Framework Analysis Results on AISHELL-1 with DeepSeek v2 and Attention Decoding
LLM Correct Error Metrics CER Change Noun Recall Cost Tokens (in/out)
sub del insert
- 4.65 0.29 0.12 5.06 - 79.32 -
+ Base 9.76 4.56 38.76 53.10 +48.04 (949%) 55.12 62k/89k
+ Pre-Detection 4.71 1.62 2.10 8.19 +3.13 (61.8%) 81.71 72k/87k
+ Chain of Thought Sub-Tasks 5.41 0.95 0.68 7.05 +1.99 (39.3%) 72.30 239k/122k
+ Iterative Correction 4.02 0.63 0.27 4.89 -0.17 (3.3%) 77.71 251k/261k
+ Answer Verification 4.00 0.49 0.20 4.69 -0.39 (7.7%) 81.05 251k/260k
IV-D Experiment Results
As detailed in Table II, GPT-4o achieved an average CER reduction of 0.89% (17.4%) on the AISHELL-1 dataset, with a maximum reduction of 1.11% (21%). As shown in Table III, GPT-4o once again demonstrated the best performance on the AISHELL-2 dataset, achieving a CER reduction of 0.62%. On the test-clean dataset, GPT-4o achieved the lowest WER of 3.19%, with an average reduction of 0.21% (5.72%) and a maximum WER reduction of 0.34% (9%). On the test-other dataset, the minimum WER was 8.32%, and the maximum relative WER reduction was 1.09% (11.4%). We also conducted experiments on Librispeech using Deepspeek v2, which resulted in bad results because Deepspeek v2 is a LLM focused on Chinese.
IV-E Result Analysis
Results show that LLMs effectively correct substitution errors in speech recognition across both Chinese and English datasets, though they may increase deletion and insertion errors generated by hallucinations. To analyze noun recall in substitution errors, we used [29] for Chinese noun filtering in AISHELL-1/2, and [30] for English noun filtering in LibriSpeech. The results indicate improved noun recall.
V Ablation Study
We analyze the different components of the RLLM-CF in V-A, demonstrate the detailed correction process of sentences in V-B on AISHELL-1 with Deepseek v2 and simply analyze the token consumption in V-C.
V-A Framework Analysis
To evaluate the performance of different components in the framework, we conduct a framework decomposition experiment on AISHELL-1 using Deepseek v2. The results are summarized in Table V. As discussed in [20], directly using LLMs with simple prompt for error correction often leads to a higher error rate as base result shows.
To address the potential redundancy and instruction execution errors in LLM output, we first introduced a pre-detection and obtained a CER of 8.19%. Subsequently, by adding the CoT subtask , the CER was further reduced to 7.05%. The main role of the CoT subtask is to constrain the LLM output space and prevent it from adding or deleting words without basis. However, LLM still cannot correctly infer the correct answer in one go. With the application of iterative reasoning in the framework, the CER was reduced to 4.80%. By implementing answer verification, the CER was further reduced to 4.67%. The main purpose of this step is to continue to reduce the errors caused by hallucinations in LLM corrections.
V-B Process Situation Example
As shown in Figure 4, the LLM ultimately identified 2,043 sentences as problematic. After multiple inferences, it maintained confidence in 1,915 sentences, abandoned 128 sentences, and the Answer Verification rejected 347 sentences. Finally, it corrected 1,568 sentences. We can observe that the LLM corrected 383 erroneous sentences to correct ones and 98 correct sentences to erroneous ones. Additionally, there were sentences with multiple errors where the number of errors was reduced, although this is more difficult to count.
Figure 4: Analysis of Results by Sentence on AISHELL-1 Using DeepSeek v2 and Attention Decode
V-C Cost Tokens
In order to save input tokens during the experiment, we usually group multiple sentences into a group for the experiment. From our limited experience, this will only slightly reduce the performance. The specific cost tokens are shown in Table V-A. Since LLM output has great randomness, the consumption of LLM output token is also not fixed, so the above value can only be used as a reference
VI Conclusion
This paper focused on the application of general LLMs to the specialized domain of speech recognition error correction. We proposed the Reliable LLM Correction Framework, which use Pre-detection and Answer Verification to mitigate the LLM hallucination problem and ensure the reliability of LLM answers. Also, We decompose the error correction task into several subtasks and incorporates iterative inference and confidence assessment. Our framework enhanced the reliability of LLMs in correcting ASR errors and we achieved better ASR result on representative datasets.