### v3

# Fewer Hallucinations, More Verification: A Three-Stage LLM-Based Framework for ASR Error Correction

Abstract
Automatic Speech Recognition (ASR) error correction aims to correct recognition errors while preserving accurate text. Although traditional approaches demonstrate moderate effectiveness, LLMs offer a paradigm that eliminates the need for training and labeled data. However, directly using LLMs will encounter hallucinations problem, which may lead to the modification of the correct text. To address this problem, we propose the Reliable LLM Correction Framework (RLLM-CF), which consists of three stages: (1) error pre-detection, (2) chain-of-thought sub-tasks iterative correction, and (3) reasoning process verification. The advantage of our method is that it does not require additional information or fine-tuning of the model, and ensures the correctness of the LLM correction under multi-pass programming. Experiments on AISHELL-1, AISHELL-2, and Librispeech show that the GPT-4o model enhanced by our framework achieves 21%, 11%, 9%, and 11.4% relative reductions in CER/WER.
I Introduction
Text error correction has become an essential component in modern speech recognition systems, aiming to reduce word error rates (WER) [lengFastCorrectFastError2021, lengFastCorrect2Fast2022, shenMaskCorrectTokens2022, lengSoftCorrectErrorCorrection2023]. Early methods typically relied on shallow neural networks or rule-based approaches to handle simple error patterns [anantaramRepairingASROutput2018, shivakumarLearningMistakesImproving2019]. Recent advances in deep learning have led to two dominant paradigms: autoregressive sequence-to-sequence models [maniASRErrorCorrection2020, dharoAutomaticCorrectionASR2016, wangASRErrorCorrection2020] and non-autoregressive edit-based models [lengFastCorrectFastError2021, lengFastCorrect2Fast2022]. Autoregressive models exploit encoder-decoder architectures with Connectionist Temporal Classification (CTC) loss [gravesConnectionistTemporalClassification2006], including translation-style correction frameworks [maniASRErrorCorrection2020, dharoAutomaticCorrectionASR2016] and entity-aware transformers [wangASRErrorCorrection2020]. In parallel, non-autoregressive edit-based models such as FastCorrect [lengFastCorrectFastError2021, lengFastCorrect2Fast2022] and SoftCorrect [lengSoftCorrectErrorCorrection2023] predict edit operations through duration modeling and integrate CTC loss with sequence-to-sequence frameworks to enhance error detection.
While conventional approaches depend heavily on large labeled datasets and task-specific training, the emergence of large language models (LLMs) offers new opportunities due to their strong contextual adaptability. LLMs have demonstrated remarkable performance across a wide range of language-related tasks, including text classification [howard2018universal], information extraction [pai2024survey], and dialogue systems [bae2022buildingAR]. This raises a fundamental question: Can a general-purpose LLM correct ASR errors directly without fine-tuning or additional information?
Current LLM applications typically follow two paradigms: prompt-based methods and fine-tuning strategies. Prompt-based methods aim to stimulate latent capabilities, employing techniques such as generative prompting for commonsense reasoning [liuGeneratedKnowledgePrompting2022] and Chain-of-Thought (CoT) prompting to enhance reasoning [weiChainthoughtPromptingElicits2023]. Fine-tuning approaches, often using low-rank adaptation (LoRA) [huLoRALowRankAdaptation2021], improve task-specific performance but remain resource-intensive. Building on the aforementioned studies, recent works have explored LLM integration into ASR error correction pipelines [ma2025asr, yang2023generative, ma2023generativelargelanguagemodels], achieving notable CER/WER reductions. However, these methods typically require either domain-specific fine-tuning or additional contextual information, which constrains their scalability in real-world applications. Therefore, leveraging the general knowledge embedded in LLMs without fine-tuning or external inputs emerges as a more practical and scalable alternative. Nevertheless, directly applying general LLMs often results in hallucination issues [minExploringIntegrationLarge2024], posing a significant challenge to reliable correction.
Figure 1: Overview of the Reliable LLM Correction Framework (RLLM-CF). The blue components represent modules executed by the LLM, while the yellow components are handled by the external program. The green block indicates the recognition result, and the black block denotes the final output. The entire error correction process follows the direction of the arrows.
To address these issues, we propose Reliable LLM Correction Framework (RLLM-CF), a novel LLM-based framework that directly leverages the pretrained knowledge of general LLMs without fine-tuning or external information. RLLM-CF mitigates hallucinations by integrating error pre-detection, reasoning-based verification, and a CoT strategy that structures correction into explainable and grounded subtasks, enabling broad applicability across domains. The proposed framework is validated on AISHELL-1, AISHELL-2, LibriSpeech test-clean and test-other datasets, achieving CER/WER reductions of 21%, 11%, 9%, and 11.4%, respectively.
The main contributions of this paper are as follows:
  • We propose RLLM-CF, an ASR error correction framework base on general-purpose LLM that requires neither fine-tuning nor external information.
  • We introduce a hallucination mitigation mechanism based on error pre-detection and reasoning-based verification, addressing the limitations of prior methods that directly accept LLM correction outputs and risk hallucination errors.
  • We employ a Chain-of-Thought (CoT) prompting strategy to enhance correction reliability by decomposing the correction process into subtasks: localization, pronunciation assessment, candidate generation, and candidate selection.
The remainder of this paper is organized as follows. Section II reviews recent LLM-based ASR correction approaches, with a particular focus on hallucination challenges. Section III details the proposed framework. Section IV presents experimental results and analysis on benchmark datasets. Section V provides further discussions, and Section VI concludes the paper.
II Related Work
II-A LLM-based ASR Error Correction
In recent years, integrating LLMs into ASR error correction pipelines has attracted increasing attention. Min and Wang [minExploringIntegrationLarge2024] investigated the direct application of LLMs for error correction and concluded that models such as GPT-4o are ineffective due to hallucination issues. Ma et al. [ma2023generativelargelanguagemodels] explored combining N-best rescoring with LLM-based correction; however, obtaining the N-best list from the ASR system may impose additional costs in practice. Yang et al. [yang2023generative] compared LLM-based rescoring and generation methods, demonstrating that the latter outperforms the former when domain-specific information is provided or the LLM is fine-tuned. Nevertheless, such approaches require either domain knowledge or fine-tuning.
In contrast to existing methods, our approach eliminates the need for supplementary resources (e.g., N-best lists or domain-specific knowledge), avoids fine-tuning the LLM, and incorporates critical verification mechanisms to ensure correction reliability without blind reliance on model outputs.
II-B A Potential Catastrophe: Hallucination Risks in LLMs
Hallucinations in LLMs refer to instances where the generated responses, while grammatically correct, fluent, and plausible, deviate from the input or contradict factual information [huang2023survey, bai2024hallucination]. In the context of ASR error correction, model-induced hallucinations are closely linked to transcription errors. Min and Wang [minExploringIntegrationLarge2024] conducted an empirical study on directly integrating LLMs into ASR systems, confirming that powerful generative models can inadvertently introduce hallucination errors by fabricating content absent from the original audio.
In this work, we further validate and analyze the hallucination phenomenon arising from direct LLM-based ASR error correction using simple prompts, as detailed in Section IV-A. Moreover, the proposed method effectively mitigates hallucinations, resulting in more reliable and accurate correction performance.
III Reliable ASR Error Correction Framework with LLM
This section presents the overall architecture of RLLM-CF, including the design of the error pre-detection module, the decomposition of ASR error correction into subtasks via a Chain-of-Thought strategy, and the implementation of an answer verification mechanism. Detailed descriptions of each component are provided in Sections III-A, III-C and III-D.
III-A Reliable LLM Correction Framework
The overall framework of RLLM-CF is illustrated in Figure 1. It consists of three main components: Error Pre-Detection, Chain-of-Thought Subtask Iterative Correction, and Reasoning Process Verification. The prompt design, illustrated in Figure III-D, incorporates three additional few-shot correction examples demonstrating specific reasoning steps, which are omitted for clarity. For Chinese error correction, we use the translated version of this prompt. The complete correction algorithm is outlined in Algorithm 1.
Figure 2: The above picture introduces the specific parts of Prompt’s design in detail.;The text in the white boxes represents the specific prompt content, while the blue boxes provide explanations.
1
2Define ;
3 $y$ = Input ASR Result;
4 $\mathcal{R}$ = LLM Correction Reasoning CoT;
5 $\mathcal{P}$ = Chain of Thought Sub-Tasks Prompt;
6 $\mathcal{G}$ = LLM Output Space Limited under Prompt;
7
8 Stage 1: Error Detection;
9 $\mathcal{D}(y)=\begin{cases}1,&\text{if LLM detects errors in }y\\ 0,&\text{otherwise}\end{cases}$;
10
11 Stage 2: Iterative Correction;
12 if $\mathcal{D}(y)==0$ then
13 return y
14 end if
15else
[…]
17 $\mathcal{Y}\leftarrow\emptyset$;
18 for $k\leftarrow 1$ to max_steps do
19    $y^{*},R\leftarrow\underset{y^{*}\in\mathcal{G}}{\arg\max}\ \mathbb{P_{LLM}}(y^{*}|y,\mathcal{P},\mathcal{Y})$;
20    if $\mathcal{C}(y^{*},y)==1$ then
21       return $y^{*},R$
22      end if
23    $\mathcal{Y}\leftarrow\mathcal{Y}\cup\{y^{*}\}$;
24
25   end for
26
27 end if
28
29 Stage 3: Answer Validation;
30 if R exists and $\mathcal{R}$ passes Answer Verification then
31 return $y^{*}$
32 end if
33else
34 return y
35 end if
Algorithm 1 Three-Stage Reliable LLM Correction Algorithm
TABLE I: LLMs Correction Hallucinations with plain Prompt: ”Please help me correct the text generated by speech recognition. Please note that only the corrected text needs to be output.” and without additional processing
Type Category Example Explanation
Faithful Hallucinations Instruction Violation Input: where is the nist hotel The LLM failed to follow the instruction
Output: Sorry, i can’t answer this question.
Redundant Output Input: i don’t know all of them but i know linden’s are The LLM produces additional outputs while correcting
Output: This answer is: i don’t know all of them but i know linden’s are.
Continue Writing Input: stephen leaning back and drawing idly on his scribbler listened to the talk about him which heron checked from time to time by saying The LLM expands the original text while correcting it
Output: Stephen leaning back and drawing idly on his scribbler listened to the talk about him which heron checked from time to time by saying ”Due to their praise of maiden pure of teeming motherhood.”
Blank Output Input: i discovered and put out a fire that would have destroyed the whole plant but marshall never even thanked me The LLM produces blank output
Output:
Repeated Output Input: reenter butler and three footmen who moved the tea things hostess to two guests The LLM produces repeated outputs
Output: Reenter butler and three footmen who moved the tea things hostess two two two two two two two two two two…
Grammar Correction Input: it must remember be one or the other The LLM corrects grammar error
Output: It must be remembered to be one or the other.
Factual Hallucinations Make Mistake Input: i believe in the trainin of people to their hask capacity the englishman here heartily seconded him The LLM makes a mistake
Output: I believe in the trainin of people to their task capacity the englishman here heartily seconded him.
III-B Error Pre-Detection
As shown in Table I, general LLMs often generate hallucinations, producing more errors than corrections. This highlights the necessity of adopting an error prevention first strategy for correction tasks. To prevent LLMs from altering correct content, we first instruct the model to detect errors in the input sentence. If no errors are detected, the sentence is directly retained; otherwise, the model proceeds to the correction stage, referred to as Stage 1 in Algorithm 1.
III-C Chain of Thought Subtask Iterative Correction
During error correction, denoted as Stage 2 in Algorithm 1, LLMs may fail to infer the correct answer or introduce hallucinations by making unwarranted modifications. To address this, we decompose the correction task into four subtasks—localization, pronunciation assessment, candidate generation, and candidate selection—following a CoT strategy to improve reasoning reliability, as illustrated in Figure 3.
Despite this structured reasoning, the generated results may still contain errors. To mitigate this, we perform a confidence evaluation on the generated answer. If the model remains confident, the corrected answer is output in the required format; otherwise, the correction process iterates until either a valid answer is obtained or the maximum number of iterations is reached.
III-D Answer Verification
Following the correction, a verification step is conducted to ensure compliance with task instructions. Specifically, we employ the model’s output to assess: (1) whether the answer conforms to the required format, and (2) whether all reasoning steps are correctly completed. Only when both criteria are satisfied is the corrected output accepted; otherwise, the original input is retained.
Figure 3: The above is a detailed answer to illustrate the composition and specific answers of the subtasks.
IV Experiment
The experimental section is organized as follows. Section IV-A investigates and validates LLM hallucination phenomena. Section IV-B and Section IV-C describe the experimental setup, including dataset selection and model configurations. Experimental results on AISHELL-1 [buAISHELL1OpensourceMandarin2017], AISHELL-2 [duAISHELL2TransformingMandarin2018], and LibriSpeech [panayotovLibrispeechASRCorpus2015] are reported in Section IV-D. Finally, Section IV-E presents an overall performance analysis, including a detailed evaluation of noun recall rates.
IV-A LLM Hallucination
In experiments with basic prompting, we categorize LLM hallucinations in ASR error correction into two types: faithful hallucinations—including instruction violations, redundant outputs, continuation of writing, blank outputs, and grammar corrections—and factual hallucinations, characterized by content errors. A detailed categorization is summarized in Table I.
IV-B The Dataset Configuration
TABLE II: Character Error Rate (CER) and Noun Recall of GPT-4o and Deepseek v2 on AISHELL-1 ASR Error Correction;The ASR model is a conformer-based AED model trained on AISHELL-1.
Decode Method LLM Correction Error Metrics Performance Metrics
sub del insert Change CER Noun Recall
Attention[chan2016listen] 4.65 0.29 0.12 - 5.06 79.32
Attention Rescore[zhang2020unified] 4.42 0.12 0.08 - 4.62 79.93
CTC Greedy Search[gravesConnectionistTemporalClassification2006] 4.95 0.13 0.09 - 5.17 77.91
CTC Prefix Search[gravesConnectionistTemporalClassification2006] 4.95 0.13 0.09 - 5.17 77.88
Attention[chan2016listen] GPT-4o 3.67 0.44 0.21 -0.74(14.6%) 4.32 82.11
Attention Rescore[zhang2020unified] 3.45 0.34 0.22 -0.61(13%) 4.01 82.96
CTC Greedy Search [gravesConnectionistTemporalClassification2006] 3.56 0.32 0.18 -1.11(21%) 4.06 82.24
CTC Prefix Search[gravesConnectionistTemporalClassification2006] 3.56 0.30 0.20 -1.11(21%) 4.06 81.95
Attention[chan2016listen] Deepseek v2 4.00 0.49 0.20 -0.37(7.3%) 4.69 81.05
Attention Rescore[zhang2020unified] 3.73 0.30 0.18 -0.41(8.8%) 4.21 81.98
CTC Greedy Search[gravesConnectionistTemporalClassification2006] 3.96 0.35 0.20 -0.66(12.7%) 4.51 80.62
CTC Prefix Search[gravesConnectionistTemporalClassification2006] 3.95 0.36 0.17 -0.69(13%) 4.48 80.46
TABLE III: Word Error Rate (WER) and Noun Recall of GPT-4o on Librispeech ASR Error Correction (test-clean/other Subsets).The ASR Model is conformer-based AED trained on Librispeech.
Decode Method LLM Correction Error Metrics Performance Metrics
sub del insert Change WER Noun Recall
Attention[chan2016listen] 2.71/6.68 0.72/1.14 0.40/0.96 - 3.82/8.79 94.76/88.20
Attention Rescore[zhang2020unified] 2.72/7.05 0.28/0.81 0.36/0.90 - 3.35/8.77 95.00/87.91
CTC Greedy Search[gravesConnectionistTemporalClassification2006] 3.00/7.59 0.33/0.99 0.37/0.93 - 3.77/9.52 94.40/86.93
CTC Prefix Search[gravesConnectionistTemporalClassification2006] 3.05/7.58 0.32/0.98 0.38/0.95 - 3.75/9.50 94.42/87.05
Attention[chan2016listen] GPT-4o 2.43/6.00 0.73/1.19 0.47/1.12 -0.18/-0.47 3.64/8.32 95.67/90.40
Attention Rescore[zhang2020unified] 2.47/6.39 0.25/0.73 0.46/1.26 -0.16/-0.39 3.19/8.38 96.05/90.10
CTC Greedy Search[gravesConnectionistTemporalClassification2006] 2.63/6.36 0.31/0.88 0.49/1.20 -0.34/-1.07 3.43/8.45 95.57/90.08
CTC Prefix Search[gravesConnectionistTemporalClassification2006] 2.75/6.30 0.32/0.89 0.51/1.22 -0.17/-1.09 3.58/8.41 95.55/89.82
TABLE IV: CER and Noun Recall of GPT-4o on AISHELL-2 Streaming ASR Error Correction. The ASR model is a conformer-based AED model trained on AISHELL-2 and decoded with the Attention Rescore Decoding Method using a chunk size of 16.
LLM Correction Error Metrics Performance Metrics
sub del insert Change CER Noun Recall
5.26 0.17 0.14 - 5.57 77.64
GPT-4o 4.47 0.29 0.19 -0.62 (11%) 4.95 79.66
We select three datasets as evaluation benchmarks, including the Chinese datasets AISHELL-1 [buAISHELL1OpensourceMandarin2017] and AISHELL-2 [duAISHELL2TransformingMandarin2018], and the English dataset LibriSpeech [panayotovLibrispeechASRCorpus2015], aiming to evaluate performance on both Chinese and English data. The LibriSpeech test set is partitioned into two subsets: test-clean, containing high-quality audio, and test-other, comprising noisier audio samples.
IV-C The Model Configuration
In our experiment, we adopt a conformer-based attention-encoder-decoder (AED) ASR model trained with the WeNet toolkit [yaoWeNetProductionOriented2021], following the U2++[wu2106u2++] method, across three datasets. For AISHELL-1 and LibriSpeech, a non-streaming model architecture is employed, and four decoding strategies are evaluated: Attention, Attention Rescoring, CTC Greedy Search, and CTC Prefix Search.
To further assess the performance of streaming models, experiments are conducted on AISHELL-2 using the Attention Rescoring decoding method with a chunk size of 16. The Conformer encoder comprises 12 blocks with an attention dimension of 256, 4 attention heads, and 2048 linear units, incorporating relative positional encoding and Swish activation. The decoder is implemented as a bidirectional Transformer with three forward and three backward layers. The network follows a hybrid CTC/attention architecture, with the CTC loss weight set to 0.3.
For LLMs, we select GPT-4o [openaiGPT4TechnicalReport2024] and DeepSeek-V2 [deepseek-aiDeepseekv2StrongEconomical2024], both demonstrating strong multilingual capabilities in Chinese and English. The decoding temperature is set to 0.2, and the top-$p$ parameter is set to 0.8.
TABLE V: Framework Analysis Results on AISHELL-1 with DeepSeek v2 and Attention Decoding Method; The ASR model is a conformer-based AED model trained on AISHELL-1
LLM Correction Error Metrics Performance Metrics Cost Tokens (in/out)
sub del insert CER Change Noun Recall
- 4.65 0.29 0.12 5.06 - 79.32 -
+ Base 9.76 4.56 38.76 53.10 +48.04 (949%) 55.12 62k/89k
+ Pre-Detection 4.71 1.62 2.10 8.19 +3.13 (61.8%) 81.71 72k/87k
+ Chain of Thought Sub-Tasks 5.41 0.95 0.68 7.05 +1.99 (39.3%) 72.30 239k/122k
+ Iterative Correction 4.02 0.63 0.27 4.89 -0.17 (3.3%) 77.71 251k/261k
+ Answer Verification 4.00 0.49 0.20 4.69 -0.39 (7.7%) 81.05 251k/260k
IV-D Experiment Results
Table II summarizes the results on the AISHELL-1 dataset, where GPT-4o achieved an average CER reduction of 0.89 percentage points (17.4% relative reduction), with a maximum reduction of 1.11 percentage points (21%). As shown in Table IV, GPT-4o achieved the better performance on the AISHELL-2 dataset, with a CER reduction of 0.62 percentage points.
As shown in Table III, on the LibriSpeech test-clean subset, GPT-4o achieved the lowest WER of 3.19%, corresponding to an average reduction of 0.21 percentage points (5.72% relative reduction) and a maximum reduction of 0.34 percentage points (9%). On the test-other subset, a WER of 8.32% was achieved, with a maximum relative reduction of 1.09 percentage points (11.4%).
Experiments with DeepSeek-V2 on LibriSpeech yielded suboptimal performance, as DeepSeek-V2 is primarily designed for Chinese and demonstrates limited capability on English datasets.
IV-E Result Analysis
The results indicate that LLMs effectively correct substitution errors in speech recognition across both Chinese and English datasets. However, a slight increase in deletion and insertion errors is observed, primarily due to hallucinations.
To analyze noun recall within substitution errors, noun filtering was performed using the method proposed in [che-etal-2021-n] for Chinese (AISHELL-1/2) and [bird2009natural] for English (LibriSpeech). The analysis reveals an improvement in noun recall following correction.
V Ablation Study
The ablation study analyzes the contribution of different components in V-A, presents a case study of the correction process in V-B on AISHELL-1 using DeepSeek-V2, and briefly analyzes token consumption in V-C.
V-A Framework Analysis
To evaluate the effectiveness of each framework component, a component-wise ablation study was conducted on AISHELL-1 using DeepSeek-V2. Results are summarized in Table V. Consistent with prior findings [minExploringIntegrationLarge2024], using LLMs with simple prompting for error correction often leads to increased error rates, as reflected by the baseline results.
To mitigate redundancy and instruction execution errors in LLM outputs, an error pre-detection module was introduced, reducing the CER to 8.19%. Incorporating CoT subtask decomposition further reduced the CER to 7.05%, as it primarily constrains the LLM output space and prevents unwarranted word insertions or deletions. However, correct inference cannot always be achieved in a single pass.
By integrating iterative reasoning into the framework, the CER was reduced to 4.89%. Finally, implementing answer verification further reduced the CER to 4.69%, effectively mitigating hallucination-induced errors in LLM corrections.
V-B Process Situation Example
As shown in Figure 4, the LLM identified 2,043 problematic sentences. Through iterative refinement, confidence was retained for 1,915 sentences, while 128 were discarded and 347 were rejected by the answer verification module. Ultimately, 1,568 sentences were successfully modified. Among these, 383 erroneous sentences were corrected, whereas 98 originally correct sentences were mistakenly altered to erroneous forms.
For sentences containing multiple errors, the overall error count was reduced, although precise quantification remains challenging.
Figure 4: Analysis of Results by Sentence on AISHELL-1 Using DeepSeek-V2 and Attention Decoding Method; The ASR model is a conformer-based AED model trained on AISHELL-1
V-C Token Consumption Analysis
To reduce input token consumption during experiments, multiple sentences were grouped together for inference. Empirical observations suggest that this strategy only slightly degrades performance. The specific token consumption is summarized in Table V.
Since LLM outputs are inherently variable, the number of output tokens also fluctuates across runs; thus, the reported values serve only as approximate references.
VI Conclusion
This paper focused on the application of general LLMs to the specialized domain of speech recognition error correction. We proposed the Reliable LLM Correction Framework, which use Pre-detection and Answer Verification to mitigate the LLM hallucination problem and ensure the reliability of LLM answers. Also, We decompose the error correction task into several subtasks and incorporates iterative inference and confidence assessment. Our framework enhanced the reliability of LLMs in correcting ASR errors and we achieved better ASR result on representative datasets.